find three consecutive integers such that the sum of the first two integers is 24 more than the third one
Answer by harpazo(655)

x + x + 1 = 24 + x + 2

2x + 1 = x + 26

2x - x = 26-1

x = 25

The three consecutive numbers are 25, 26, and 27.

Did you follow?